# Elevated UV luminosity density at Cosmic Dawn explained by non-evolving, weakly-mass dependent star formation efficiency

### v1

@@ Appendix B Masses of halos and galaxies
Figure 13: Halo mass (left panels) and stellar mass (right panels) in bins of UV magnitude for galaxies at $z\sim{}6-15$ in FIREboxHR. The larger panels at the top show the logarithm of the average mass in each UV magnitude bin while the smaller panels at the bottom show the corresponding scatter of the mass in dex. Filled symbols and thick lines show results for FIREboxHR for different redshifts (see legend). Bins with less than 3 data points are not shown. The figure only includes halos with at least 300 dark matter particles ($M_{\rm halo}\gtrsim{}10^{7}\,M_{\odot}$). UV magnitudes shown on the x-axis include the effects of dust attenuation. Uncertainties (16th–84th percentiles) are computed via bootstrapping. Empty squares and diamonds in the top left panel are observational estimates at $z\sim{}6-7$ based on galaxy clustering by Harikane et al. (2022). Thin lines show the observed $M_{\rm UV}-M_{\rm star}$ relation at $z\sim{}6-10$ (Song et al., 2016; Stefanon et al., 2021; Endsley et al., 2023a). The normalization of this relation is sensitive to the adopted star formation history, see, e.g., Endsley et al. (2023a). On average, galaxies that are more UV luminous are more massive and reside in more massive halos. Stellar and halo masses at fixed UV magnitude increase with decreasing redshift. The scatter ($\sim{}0.3-0.4$ dex) in either relation is significant.
@@
The evolution of the average halo and stellar mass at fixed UV luminosity can be understood in the context of a non-evolving SFE – halo mass relation. A galaxy in a halo of any given mass will, on average, have a higher UV luminosity at a higher $z$ because the specific halo growth rate increases with $z$ at fixed mass (Behroozi & Silk, 2015) and the UV luminosity is proportional to the halo growth rate and the SFE (eq. 9). This observation combined with a comparably low and approximately mass and redshift independent scatter implies that the average halo mass hosting galaxies of a given UV luminosity will decrease with increasing redshift. Given that the stellar mass – halo mass relation has low scatter and is approximately redshift independent at high $z$ in FIRE-2 zoom-in simulation (Ma et al., 2018, 2019), the stellar mass will also decrease with increasing redshift at fixed UV magnitude.
@@ Appendix C SFE for different averaging times of star formation
Appendix D Fitting parameters for the UV luminosity functions

### v2

@@ Appendix B Masses of halos and galaxies
Figure 13: Halo mass (left panels) and stellar mass (right panels) in bins of UV magnitude for galaxies at $z\sim{}6-15$ in FIREboxHR. The larger panels at the top show the logarithm of the average mass in each UV magnitude bin while the smaller panels at the bottom show the corresponding scatter of the mass in dex. Filled symbols and thick lines show results for FIREboxHR for different redshifts (see legend). Bins with less than 3 data points are not shown. The figure only includes halos with at least 300 dark matter particles ($M_{\rm halo}\gtrsim{}10^{7}\,M_{\odot}$). UV magnitudes shown on the x-axis include the effects of dust attenuation. Uncertainties (16th–84th percentiles) are computed via bootstrapping. Empty squares and diamonds in the top left panel are observational estimates at $z\sim{}6-7$ based on galaxy clustering by Harikane et al. (2022). Thin lines show the observed $M_{\rm UV}-M_{\rm star}$ relation at $z\sim{}6-10$ (Song et al., 2016; Stefanon et al., 2021; Endsley et al., 2024). The normalization of this relation is sensitive to the adopted star formation history, see, e.g., Endsley et al. (2024). On average, galaxies that are more UV luminous are more massive and reside in more massive halos. Stellar and halo masses at fixed UV magnitude increase with decreasing redshift. The scatter ($\sim{}0.3-0.4$ dex) in either relation is significant.
@@
The evolution of the average halo and stellar mass at fixed UV luminosity can be understood in the context of a non-evolving SFE – halo mass relation. A galaxy in a halo of any given mass will, on average, have a higher UV luminosity at a higher $z$ because the specific halo growth rate increases with $z$ at fixed mass (Behroozi & Silk, 2015; Rodríguez-Puebla et al., 2016) and the UV luminosity is proportional to the halo growth rate and the SFE (Eq. 9). This observation combined with a comparably low and approximately mass and redshift independent scatter implies that the average halo mass hosting galaxies of a given UV luminosity will decrease with increasing redshift. Given that the stellar mass – halo mass relation has low scatter and is approximately redshift independent at high $z$ in FIRE-2 zoom-in simulation (Ma et al., 2018, 2019), the stellar mass will also decrease with increasing redshift at fixed UV magnitude.
@@ Appendix C SFE for different averaging times of star formation
Appendix D The faint end slope of the UV luminosity function
Eq. (16) expresses the UV luminosity function (UV LF) in terms of the halo mass function (HMF) $dn/dM$ and the probability distribution $p(\mathcal{M}=m|{}M,z)$ of the UV magnitude $\mathcal{M}$ as function of halo mass $M$ and redshift $z$. Following the discussion in Section 3.4, we model $p(\mathcal{M}=m|{}M,z)$ as a normal distribution with the mean $\langle{}\mathcal{M}\rangle{}$ and standard deviation $\sigma_{\mathcal{M}}$ given by equation (12). The UV luminosity function can thus be written as
[…]
with $\mathscr{M}={\rm Mag}(\langle{}\mathcal{S}\rangle\dot{M}/\kappa)$ and $c=\ln(10)/2.5$.
We can solve Eq. (23) analytically if the halo mass function, the SFE, and the halo growth rate are all power-law functions of the halo mass, and $\sigma_{\mathcal{M}}$ is a constant, i.e., if
$\frac{dn}{dM}$ $=\left(\frac{dn}{dM}\right)_{0}\,\tilde{M}^{\alpha_{M}},$ $\langle{}\mathcal{S}\rangle$ $=\langle{}\mathcal{S}\rangle_{0}\,\tilde{M}^{\beta},$ $\dot{M}$ $=\dot{M}_{0}\,\tilde{M}^{\gamma},$ (24)
where $\tilde{M}=M/M_{0}$, $M_{0}$ is a chosen reference halo mass, and pre-factors with the subscript 00 do not depend on halo mass but potentially on other quantities such as redshift. Inserting Eq. (24) into Eq. (23) and integrating over $M$ results in
$\phi(m,z)=c\phi_{*}10^{-0.4(1+\alpha)(m-m_{*})},$ (25)
with
$\alpha$ $=\frac{1+\alpha_{M}}{\beta+\gamma}-1$ (26)
$m_{*}$ $={\rm Mag}(\langle{}\mathcal{S}\rangle_{0}\dot{M}_{0}/\kappa)$ (27)
$\phi_{*}$ $=\left(\frac{dn}{dM}\right)_{0}M_{0}\frac{1}{\beta+\gamma}\exp\left(\frac{c^{2}}{2}(1+\alpha)(2+\alpha)\sigma_{\mathcal{M}}^{2}\right).$ (28)
According to Eq. (25), the UV LF is then a power law with respect to UV luminosities. Moreover, since Eq. (25) has the same functional form as the faint magnitude limit of a traditional Schechter or double power law function, we can identify the following parameters: $\alpha$ represents the faint-end slope, $m_{*}$ corresponds to the characteristic magnitude, and $\phi_{*}$ is the normalization of the luminosity function. Due to the scale-free nature of a power law, either $m_{*}$ or $\phi_{*}$ (but not both independently) can be freely adjusted by selecting an appropriate value for $M_{0}$. Multiplying the normalization $\langle{}\mathcal{S}\rangle_{0}$ of the SFE – halo mass relation by a factor $f$ increases $m_{*}$ by $-2.5\lg{}f$, i.e., it amounts to a horizontal shift by this amount. Alternatively, such a shift may be interpreted as a vertical increase of $\lg{}\phi$ by $-(1+\alpha)\lg{}f$.
Eq. (26) shows that the faint-end slope $\alpha$ of the UV LF is directly related to the combination of $\alpha_{M}$ and $\beta+\gamma$. Provided the scatter $\sigma_{M}$ is mass independent as assumed in the derivation above or small enough to be ignored, we can directly infer the low-mass slope of the SFE–halo mass relation (denoted by $\beta$ here and by $\alpha_{1}$ in the main text) from measurements of the faint-end slope of the UV LF for a given cosmological model with well constrained $\alpha_{M}$ and $\gamma$. Alternatively, if $\beta$ can be measured independently by other means, the faint end of the UV LF may be used to constrain the growth properties of halos ($\alpha_{M}$ and $\gamma$) and perhaps even test alternative dark matter models.
The faint-end slope also determines how a larger scatter in the SFE, given by ${\rm std}(\lg{}\mathcal{S})=\sigma_{\mathcal{M}}/2.5$, affects the normalization of the faint end of the UV LF. According to Eq. (28), an increase in the scatter of the SFE will raise the normalization of the UV LF when $\alpha<-2$ or $\alpha>-1$ and decrease $\phi_{*}$ when $-2<\alpha<-1$. These considerations clarify how the UV LFs shown in the right panel of Fig. 7 are affected by the scatter in the SFE. At sufficiently faint magnitudes, a power-law approximation of the UV LF results in a slope $\alpha$ that is shallower than $-2$ because of the modest turn-over at faint magnitudes seen in FIREboxHR. As a result, a larger scatter reduces the UV LF at the faintest magnitudes. In contrast, at brighter UV magnitudes, the UV LF has a local slope steeper than $-2$, leading to an increase in number density with increasing scatter. The scatter has little impact on the UV LF when the local slope is near $-2$. This condition is met around $M_{\rm UV}\sim{}-18$ at $z=6$, and $M_{\rm UV}\sim{}-15$ at $z=14$ for the model with a fixed scatter of 0.4 dex shown in Fig. 7. This result explains why the UV luminosity density integrated down to $M_{\rm UV}=-17$ is only weakly dependent on the scatter of the SFE as shown in the left panel of Fig. 9.
Appendix E Fitting parameters for the UV luminosity functions